A positive integer has 1001 digits all of which are 1 if this integer is divided by 1001 then find remainder

Grade:

1 Answers

Srutarshi Tripathi
3 years ago
The number N: (101002 – 1)/9 
Now, by Fermat’s Little Theorem , 101001 = 10 (mod 1001)
or, 101002 – 1 = 99 (mod 1001)
or, (101002 – 1)/9  = 11 (mod 1001)
 
Hence, THE REQUIRED REMAINDER IS 11 (ANSWER)